# Efficient Observation Time Window Segmentation for Administrative Data Machine Learning

### 29 Jan 2024 v1

@@ 4.3. Feature Set Creation
DI MIMIC $L$ $V$ Percent Empty $V$ Percent Empty 1 23 76.26 48 76.182 90 2070 86.10 4320 89.291 Table 3. Size and sparsity of $D_{\text{All}}$
DI MIMIC $L$ $V$ Percent Empty $V$ Percent Empty 1 23 76.02 48 76.07 90 290 86.10 315 78.87 Table 4. Size and sparsity of $D_{3}$
4.4. Algorithms
@@ 4.4.1. Classic Models
Logistic Regression: Logistic regression is selected as a simple baseline algorithm and due to its popularity for healthcare applications (cox1958, 5). The version of logistic regression used in this paper is from Sci-kit Learn with class weights inversely proportional to class frequencies due to the imbalance nature of the data (scikit-learn-logistic-regression, 36). The solver defaults to the limited memory variant of the BFGS algorithm with a regularization strength of 1.0.
@@
Random Forest: Random Forest is selected since it is robust against overfitting and adept at handling a large number of features (groll2018prediction, 9). Implementation utilizes Sci-kit Learn with class weights inversely proportional to class frequencies (scikit-learn-random-forest, 35). The number of trees in the forest is 300 and the maximum number of features is set to the root of all the features submitted to the model. The minimum samples per leaf is set to 1 and the minimum samples per split is 2. The max depth per tree was set to 30.
4.4.2. Deep Learning Models
@@
Deep Learning models leverage neural networks to extract complex patterns from large data sets (LeCun2015, 19) and are considered due to the superior performance they offer for a variety of applications. In this study, the DL models employ the ReLU activation (Glorot2011, 7) function for hidden layers to mitigate the vanishing gradient problem and a sigmoid (Goodfellow-et-al-2016, 8) function in the final layer for binary classification. The Adam optimizer (Kingma2014, 15) with varying learning rates is used for its efficiency in large-scale applications. Dropout regularization is applied to each dense layer (except the final one) to prevent overfitting, with dropout (Srivastava2014, 37) rates varying between 0.1 and 0.4. The models use binary cross-entropy as the loss function and mini-batch gradient descent with different batch sizes for optimization (Goodfellow-et-al-2016, 8).
@@
Convolutional Neural Networks (CNNs): CNNs, traditionally used in image processing, have been adapted for time-series data analysis (kashiparekh2019convtimenet, 14). As shown in Table 5, the CNN utilized in this paper has 1D convolutional layers, followed by batch normalization for stabilized learning, and max pooling layers for dimensionality reduction (Goodfellow-et-al-2016, 8) . This approach allows our CNN model to capture both spatial and temporal patterns in the data. The flattened data is then processed by dense layers, combining spatial hierarchies with classification tasks which makes CNNs particularly effective for analyzing the different temporal time patterns present in the administrative data (Michelucci2019AdvancedAD, 26).
@@
Multi-layer Perceptrons (MLPs): Multi-layer perceptrons (MLPs) are a fundamental type of neural network made of dense layers that excel at discerning complex patterns within input data (Goodfellow-et-al-2016, 8). Their relative simplicity over more complex deep learning structures are useful in interpreting the results presented in Section 5. Table 5 presents three different MLP configurations for a low (small), medium (average) and high (large) complexity model which represent different compromises between performance and the resources required for training (Calin2020DeepLA, 1). The effect of MLP model complexity on the impact of time bin size optimization will be demonstrated in Section 5.
@@
Gated Recurrent Units (GRUs): GRUs are a streamlined variant of recurrent neural networks (RNNs) that are particularly adept at processing sequential and temporal data. Their architecture is simpler than long short-term memory (LSTM) networks and involves two gates - reset and update - that efficiently modulate the flow of information (cho-etal-2014-properties, 4). This structure not only allows for more efficient computation but also effectively captures temporal dependencies, crucial for learning from sequences. GRUs have been shown to be particularly effective for administrative data (che2018recurrent, 3). The layered structured of the GRU model used in this paper is shown in Table 5.
@@
Table 5. Deep learning model parameters.
5. Results
5.1. Framework of the Experiments
@@
5.2. Experiment 1: Model Class and Architecture Variations
@@
Figure 2. Experiment 1: Chronic shelter use (DI data set).
Figure 3. Experiment 1: Patient mortality (MIMIC-III data set).
@@
5.3. Experiment 2: Consistent Model Class and Architecture
@@
Feature Scores Sleep (1Hot) 1.8495 ProgressDetails (1Hot) 0.3895 EMS 0.3688 Log (1Hot) 0.3584 Bar (1Hot) 0.0651 Table 6. Chronic Clients Feature Scores Blood Pressure 0.8842 Heart Rate 0.8833 Cancer 0.8268 Temperature 0.5667 GSC_Motor_1 (1Hot) 0.5067 Table 7. Mortality Prediction
@@
Figure 4. Experiment 2: Chronic shelter use (DI data set).
Figure 5. Experiment 2: Patient mortality (MIMIC-III data set).
5.4. Experiment 3: Different Model Classes and Architectures
@@
Figure 6. Experiment 3: Chronic shelter use (DI data set).
Figure 7. Experiment 3: Patient mortality (MIMIC-III data set).
6. Discussion
6.1. Time Bin Size Importance Depends on the Problem
@@
Feature Scores Sleep (1Hot) 0.03659 Police/Justice 0.03461 Age 0.01260 ProgressDetails (1Hot) 0.0081 EMS 0.00595 Table 8. Chronic Clients Feature Scores Age 0.1238 AdmitType_2 (1Hot) 0.0710 ServiceType_3 (1Hot) 0.0700 HeartRate 0.0579 AdmitType_0 (1Hot) 0.0442 Table 9. Mortality Prediction
6.2. Selective Use of Time Bins Helps Efficiency and Performance
7. Conclusion
7.1. Summary
This paper presents the first systematic framework for optimizing observation window time bin size for administrative data machine learning applications. Using administrative data from the healthcare and housing/homelessness domains, our results show that using the TAIB algorithm to properly identify which features benefit from a higher time resolution will greatly simplify machine learning model training and can result in improved performance. By considering two different classification problems and two different data sets, this paper also demonstrates that not all problems will benefit the same amount from optimizing time bin size. Variability between administrative data machine learning problems is another reason why a systematic method for optimizing time bin size is important.
7.2. Limitations and Future Possibilities
Our research approached windowing as a binary decision, where features with a lower cluster separation score either had windowing applied or not. Future work can explore assigning optimal values of $L_{k}$ to each feature, allowing for a more tailored approach to windowing. Additionally, extending the methodology to other predictive tasks like regression, multi-class classification, anomaly detection, and clustering could offer broader insights.
Future research could also explore how TAIB could be adapted to multi-dimensional data and integrating alternative distance measures like Mahalanobis (Mahalanobis2008, 24) and correlation-based distances. This could potentially refine the algorithm’s performance and offer new perspectives on handling complex, multidimensional datasets. TAIB could also be combined with a pre-processing method that assesses a feature’s predictive power. This may lead to a single method that will select a feature if it both benefits from a higher time resolution and has sufficient predictive power to influence the overall performance of the machine learning model when its time resolution is optimized.

### 12 Mar 2024 v2

@@ 4.3 Feature Set Creation
DI MIMIC $L$ $V$ Percent Empty $V$ Percent Empty 1 23 76.26 48 76.182 90 2070 86.10 4320 89.291 Table 3: Size and sparsity of $D_{\text{All}}$
DI MIMIC $L$ $V$ Percent Empty $V$ Percent Empty 1 23 76.02 48 76.07 90 290 86.10 315 78.87 Table 4: Size and sparsity of $D_{3}$
4.4 Algorithms
@@ 4.4.1 Classic Models
Logistic Regression: Logistic regression is selected as a simple baseline algorithm and due to its popularity for healthcare applications [5]. The version of logistic regression used in this paper is from Sci-kit Learn with class weights inversely proportional to class frequencies due to the imbalance nature of the data [34]. The solver defaults to the limited memory variant of the BFGS algorithm with a regularization strength of 1.0.
@@
Random Forest: Random Forest is selected since it is robust against overfitting and adept at handling a large number of features [9]. Implementation utilizes Sci-kit Learn with class weights inversely proportional to class frequencies [33]. The number of trees in the forest is 300 and the maximum number of features is set to the root of all the features submitted to the model. The minimum samples per leaf is set to 1 and the minimum samples per split is 2. The max depth per tree was set to 30.
4.4.2 Deep Learning Models
@@
Deep Learning models leverage neural networks to extract complex patterns from large data sets [18] and are considered due to the superior performance they offer for a variety of applications. In this study, the DL models employ the ReLU activation [7] function for hidden layers to mitigate the vanishing gradient problem and a sigmoid [8] function in the final layer for binary classification. The Adam optimizer [15] with varying learning rates is used for its efficiency in large-scale applications. Dropout regularization is applied to each dense layer (except the final one) to prevent overfitting, with dropout [35] rates varying between 0.1 and 0.4. The models use binary cross-entropy as the loss function and mini-batch gradient descent with different batch sizes for optimization [8].
@@
Convolutional Neural Networks (CNNs): CNNs, traditionally used in image processing, have been adapted for time-series data analysis [14]. As shown in Table 5, the CNN utilized in this paper has 1D convolutional layers, followed by batch normalization for stabilized learning, and max pooling layers for dimensionality reduction [8] . This approach allows our CNN model to capture both spatial and temporal patterns in the data. The flattened data is then processed by dense layers, combining spatial hierarchies with classification tasks which makes CNNs particularly effective for analyzing the different temporal time patterns present in the administrative data [25].
@@
Multi-layer Perceptrons (MLPs): Multi-layer perceptrons (MLPs) are a fundamental type of neural network made of dense layers that excel at discerning complex patterns within input data [8]. Their relative simplicity over more complex deep learning structures are useful in interpreting the results presented in Section 5. Table 5 presents three different MLP configurations for a low (small), medium (average) and high (large) complexity model which represent different compromises between performance and the resources required for training [2]. The effect of MLP model complexity on the impact of time bin size optimization will be demonstrated in Section 5.
@@
Gated Recurrent Units (GRUs): GRUs are a streamlined variant of RNNs that are particularly adept at processing sequential and temporal data. Their architecture is simpler than long short-term memory (LSTM) networks and involves two gates - reset and update - that efficiently modulate the flow of information [4]. This structure not only allows for more efficient computation but also effectively captures temporal dependencies, crucial for learning from sequences. GRUs have been shown to be particularly effective for administrative data [3]. In this paper, we utilize a traditional GRU structure that utilizes data in temporal matrix form. The layered structured of the GRU model used in this paper is shown in Table 5.
@@
Table 5: Deep learning model parameters.
5 Results
5.1 Framework of the Experiments
@@
5.2 Experiment 1: Model Class and Architecture Variations
@@
Figure 2: Experiment 1: Chronic shelter use (DI data set).
Figure 3: Experiment 1: Patient mortality (MIMIC-III data set).
@@
5.3 Experiment 2: Consistent Model Class and Architecture
@@
Feature Scores Sleep (1Hot) 1.8495 ProgressDetails (1Hot) 0.3895 EMS 0.3688 Log (1Hot) 0.3584 Bar (1Hot) 0.0651 Table 6: Chronic Clients Feature Scores Blood Pressure 0.8842 Heart Rate 0.8833 Cancer 0.8268 Temperature 0.5667 GSC_Motor_1 (1Hot) 0.5067 Table 7: Mortality Prediction
@@
Figure 4: Experiment 2: Chronic shelter use (DI data set).
Figure 5: Experiment 2: Patient mortality (MIMIC-III data set).
5.4 Experiment 3: Different Model Classes and Architectures
@@
Figure 6: Experiment 3: Chronic shelter use (DI data set).
Figure 7: Experiment 3: Patient mortality (MIMIC-III data set).
6 Discussion
6.1 Assigning Different Time Bins to Features Helps Efficiency and Performance
6.2 Time Bin Size Importance Depends on the Problem
@@
Feature Scores Sleep (1Hot) 0.03659 Police/Justice 0.03461 Age 0.01260 ProgressDetails (1Hot) 0.0081 EMS 0.00595 Table 8: Chronic Clients Feature Scores Age 0.1238 AdmitType_2 (1Hot) 0.0710 ServiceType_3 (1Hot) 0.0700 HeartRate 0.0579 AdmitType_0 (1Hot) 0.0442 Table 9: Mortality Prediction
7 Conclusion
7.1 Summary
This paper presents the first systematic framework for reducing the observation window time bin size hyperparameter search space that can be applied to general administrative data. Using administrative data from the healthcare and housing/homelessness domains, our results show that using the TAIB algorithm to properly identify which features benefit from a higher time resolution will greatly simplify machine learning model training and can result in improved performance. By considering two different classification problems and two different data sets, this paper also demonstrates that not all problems will benefit the same amount from optimizing time bin size. Variability between administrative data machine learning problems is another reason why a systematic method for optimizing time bin size is important.
7.2 Limitations and Open Problems
Our research approached windowing as a binary decision, where features with a lower cluster separation score either had windowing applied or not. Future work can explore assigning optimal values of $L_{k}$ to each feature, allowing for a more tailored approach to windowing. In particular, features with higher TAIB scores can be tuned using a large number of possible time bin sizes and the number of possible bins sizes could be gradually reduced for features with lower scores.
Future research could also explore how TAIB could be adapted to multi-dimensional data and integrating alternative distance measures like Mahalanobis [23] and correlation-based distances. This could potentially refine the algorithm’s performance and offer new perspectives on handling complex, multidimensional datasets. TAIB could also be combined with a pre-processing method that assesses a feature’s predictive power. This may lead to a single method that will select a feature if it both benefits from a higher time resolution and has sufficient predictive power to influence the overall performance of the machine learning model when its time resolution is optimized.
@@ 7.2 Limitations and Open Problems
8 Acknowledgments
The authors would like to acknowledge the support of Making the Shift, the Calgary Drop-In Centre and the Government of Alberta. This study is based in part on data provided by Alberta Seniors, Community and Social Services. The interpretation and conclusions contained herein are those of the researchers and do not necessarily represent the views of the Government of Alberta. Neither the Government of Alberta nor Alberta Seniors, Community and Social Services express any opinion related to this study.